### v1

# PlanQA: A Benchmark for Spatial Reasoning in LLMs using Structured Representations

Abstract
We introduce PlanQA, a diagnostic benchmark for evaluating geometric and spatial reasoning in large-language models (LLMs). PlanQA is grounded in structured representations of indoor scenes, such as kitchens, living rooms, and bedrooms, encoded in a symbolic format (e.g., JSON, XML layouts). The benchmark includes diverse question types that test not only metric and topological reasoning (e.g., distance, visibility, shortest paths) but also interior design constraints such as affordance, clearance, balance, and usability. Our results across a variety of frontier open-source and commercial LLMs show that while models may succeed in shallow queries, they often fail to simulate physical constraints, preserve spatial coherence, or generalize under layout perturbation. PlanQA uncovers a clear blind spot in today’s LLMs: they don’t consistently reason about real-world layouts. We hope that this benchmark inspires new work on language models that can accurately infer and manipulate spatial and geometric properties in practical settings.
1 Introduction
Recent progress in large language models (LLMs) has revealed strong capabilities in structured reasoning, yet spatial inference over physically grounded environments remains poorly understood. In numerous real-world applications, including architectural design, assistive planning, and embodied interaction, spatial understanding is handled through structured formats such as JSON, in which objects are specified by position, size, and orientation, rather than through images or natural language. Reasoning in these contexts requires geometric inference over symbolic layouts, not pixel-level perception.
We introduce PlanQA, a benchmark to evaluate spatial reasoning in LLMs using 2D floor plans represented in structured text-based formats. Each instance consists of a JSON-encoded layout paired with natural language questions that require the model to compute distances, evaluate placement feasibility, assess visibility, and reason about spatial constraints. PlanQA isolates symbolic spatial reasoning over inputs that mirror the abstractions used by designers, architects, and agents operating in structured environments.
Although LLMs can increasingly be used in tool-assisted pipelines, for example to invoke spatial solvers or generate code, this work focuses on models’ direct, unaided reasoning capabilities. PlanQA is designed to probe what LLMs can infer from structured input alone, without relying on external computation or visual grounding.
Specifically, our contributions are as follows:
  • We introduce a dataset of 1,800 structured 2D layouts, 600 each from kitchens, living rooms, and bedrooms, represented in JSON and paired with spatial reasoning questions.
  • We provide a diverse suite of 18,000 questions spanning geometric, functional, and design-aware reasoning, with one question per layout across multiple categories.
  • We establish structured evaluation protocols and scoring metrics that enable a fine-grained diagnosis of reasoning performance by task type and error mode.
  • We conduct a comparative analysis of 11 LLMs, including GPT-4.1 OpenAI et al. (2024) and open models such as DeepSeek-R1 DeepSeek-AI (2025) and Llama-3.3-70B-Instruct Grattafiori et al. (2024), revealing consistent failure patterns in spatial inference from symbolic input.
PlanQA provides a diagnostic substrate for measuring spatial competence in language models, with applications in layout validation, agent planning, and symbolic grounding.
2 Related Work
Prior benchmarks have explored spatial reasoning across vision and language domains. CLEVR Johnson et al. (2017) is a synthetic visual question answering dataset designed to test compositional reasoning, including basic spatial relationships. In real-world settings, SpatialSense Yang et al. (2019) focuses on recognizing spatial relations in images through adversarially mined examples. Benchmarks like BabyAI Chevalier-Boisvert et al. (2019), ALFRED Shridhar et al. (2020), and Room-to-Room (R2R) Anderson et al. (2018) integrate spatial understanding into embodied tasks, requiring agents to follow instructions involving navigation and object manipulation in simulated environments. Recent datasets such as ScanQA Azuma et al. (2022) and 3DSRBench Ma et al. (2024a) extend spatial reasoning evaluation into 3D environments, emphasizing the need for models to comprehend and reason about spatial relationships in three dimensions.
Vision-language models have advanced spatial reasoning but often handle it qualitatively. The VQA dataset Antol et al. (2015) challenges models to answer questions about images, while VL-T5 Cho et al. (2021) unifies vision-and-language tasks via text generation. Recent work on 3D scene graphs Armeni et al. (2019) introduces structured representations of environments, facilitating spatial reasoning. However, these approaches may miss fine-grained geometric details necessary for precise spatial inference. Efforts like SpatialVLM Chen et al. (2024) aim to endow vision-language models with enhanced spatial reasoning capabilities, addressing some of these limitations.
Advancements in generative models have also contributed to spatial reasoning tasks. LayoutGPT Feng et al. (2023) leverages large language models for compositional visual planning and layout generation, while Holodeck Yang et al. (2024) enables language-guided generation of 3D embodied AI environments. Similarly, AnyHome Fu et al. (2024) focuses on open-vocabulary generation of structured and textured 3D homes, highlighting the integration of language and spatial understanding in generative contexts. Infinigen Indoors (Raistrick et al., 2024) offers richly rendered 3D scenes but suffers from poor object placement, often yielding physically implausible or random layouts due to non-convergent simulated annealing. LayoutVLM Sun et al. (2024) and FirePlace Huang et al. (2025) improve layout generation via optimization and constraint solving, respectively. But they assess spatial reasoning through output realism, not direct inference. In contrast, our benchmark tests symbolic reasoning without tool-assisted refinement.
Evaluations of large language models’ spatial understanding have been conducted in studies like Evaluating Spatial Understanding of Large Language Models Yamada et al. (2024), which assesses the spatial reasoning capabilities of LLMs through structured tasks. Additionally, benchmarks such as BALROG Paglieri et al. (2025) test agentic reasoning in game environments, further exploring the spatial and decision-making abilities of language and vision-language models. While these efforts reveal important limitations in high-level spatial understanding, our benchmark isolates low-level geometric reasoning in structured layouts, providing fine-grained and task-specific insights into models’ spatial competence. Recent 3D-LLM surveys such as Ma et al. (2024b) cover tasks like navigation and interaction, but not symbolic spatial reasoning. PlanQA fills this gap by testing raw spatial competence from structured layouts without multimodal input.
PlanQA addresses the gap in existing benchmarks by directly evaluating structured spatial inference from symbolic room layouts. Unlike prior benchmarks relying on raw images or focusing on commonsense spatial language, PlanQA provides explicit spatial representations (object coordinates and dimensions) and tests models’ abilities to perform precise spatial reasoning tasks, such as calculating distances, assessing visibility, and verifying object fit within a controlled setting.
3 Method
3.1 Synthetic Layout Generation
Our initial aim was to utilize publicly available real-world floorplan datasets. However, a comprehensive review revealed critical limitations. Prominent datasets such as SUNCG (Song et al., 2017) and HouseExpo (Li et al., 2019) were withdrawn due to unresolved copyright claims, rendering them unusable. Other large-scale resources—including 3D-FRONT (Fu et al., 2020), Structured3D (Zheng et al., 2020), and InteriorNet (Li et al., 2018)—are procedurally generated but impose constraints on layout diversity, furniture semantics, or downstream reuse. Datasets like CubiCasa5K (Kalervo et al., 2019) and Rent3D (Liu et al., 2015) offer fixed architectural plans from real environments but lack furnishing annotations. RPLAN (Wu et al., 2019), despite its scale, is not publicly released, and the dataset of Di et al. (2020), while large, appears procedurally generated albeit with realtor supervision and imposes restrictions on redistribution and derived works. Given these legal, practical, and methodological constraints, we found synthetic data generation to be the most viable alternative.
We generated 1,800 synthetic indoor layouts using Gemini 2.5 Pro, a large language model fine-tuned for spatial reasoning (Team et al., 2025). While employing an LLM for both data generation and evaluation may raise concerns of circularity, our evaluation pipeline is entirely independent: the LLM is used solely for data generation, and evaluation is performed via deterministic spatial reasoning tasks.
The generation process comprises two stages. First, we specify room geometries using explicit constraints on shape, adjacency, and design principles related to circulation, symmetry, and zoning. These constraints are encoded directly in the LLM prompt. Second, each room is furnished according to style-specific rules, also defined in structured prompts, to encourage both visual realism and functional plausibility. Approximately one-third of candidate layouts are filtered out by automated spatial validity checks, which remove configurations with inaccessible furniture or implausible adjacencies. This curation pipeline yields a diverse and geometrically coherent dataset. Full prompts, generation templates, and validation scripts are provided in the supplemental materials. Representative examples of layout types are shown in Fig. 1. A breakdown of room types and layout geometries is provided in Table 1.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 1: Representative layouts from the PlanQA dataset, spanning three room types (kitchen, living room, bedroom) and three geometric configurations (rectangular, L-shaped, open). Layouts are procedurally generated using explicit spatial constraints and validated for functional feasibility.
Table 1: Layout distribution in PlanQA by room type, style, and geometry, with room-wise totals.
Room Type Style Rectangular L-Shaped Open Total
Kitchen L-Shaped (incl. counter) 208 36 105 600
U-Shaped / G-Shaped 42 128 2
Island-Based 2 0 33
Wall / Galley (compact layouts) 17 10 17
Living Room Fireplace-Centric 55 1 34 600
Conversational 66 3 37
Multi-Zone 73 176 46
TV-Focused 78 3 28
Bedroom Bedroom with Workstation 98 35 28 600
Traditional Bedroom 114 59 57
Efficient Small Bedroom 65 3 8
Welcoming Guest Bedroom 75 30 28
Total 893 484 423 1800
3.2 Question Taxonomy, Prompting, and Cognitive Scope
PlanQA assesses spatial reasoning by presenting models with a single natural language question per symbolic layout. Questions span a range of geometric, topological, and functional types, including numeric computations (distances, areas), spatial feasibility (object fit, placement), occlusion and visibility, and environment-level assessments such as missing elements or requirement violations. Examples include queries like whether a 2 m $\times$ 3 m table can be placed without blocking passage, whether the television is visible from a particular seat, or which essential appliances are absent from a proposed kitchen. Some questions require local metric reasoning; others test whether a model can simulate constraints that would arise in physical occupancy or usability. A categorized list of question types is shown in Table 2.
Each question is generated by instantiating a parameterized template against the specific geometry of a layout. These templates are natural language prompts that fill in object names, measurements, and task-specific context. Prompts are issued in zero-shot settings with no few-shot scaffolding or role-based instruction, but include consistent structural cues to support stepwise inference. Each prompt ends with the string ”Reason step by step to solve the problem.” followed by “Final Answer:”, enabling unambiguous output extraction. For instance, a placement query might read: “Given the room layout below, compute the area of the largest rectangle that can be placed in the kitchen without overlapping any object. Reason step by step to solve the problem. Final Answer: [value in square meters].”
Table 2: PlanQA question taxonomy. The example question shown is an instantiation of the template used to generate all questions of that type. Each task is labeled with a format code, cognitive skill category (see Tables 4 and 4), and the room types in which it is applied: K = kitchen, L = living room, B = bedroom.
Category Example Question Format Target Rooms
Distance What is the distance between the refrigerator and the stove? N M K, L, B
Area What is the total sitting area? N M K, L, B
Free Space What is the unoccupied floor area in the room? N T K, L, B
View Angle What is the angle between the TV’s facing direction and the viewer’s line of sight from the armchair? N V K, L, B
Repositioning How far can the ottoman be moved to the left direction until it touches another object or the wall? N G K, L, B
Max Box What is the area of the largest axis-aligned rectangle that fits in the free space? N G K, L, B
Fit/Placement Can a 2m × 3m desk table fit in the room without overlap? B G K, L, B
Shortest Path What is the valid shortest path from the stove to the exit with clearance? S G, R K, L, B
Missing Object What is a possible placement for the missing sink? S R, T K, L, B
Obstruction What are all the objects that intersect the straight line from the fireplace to the door? L V K, L, B
Table 3: Answer format and scoring codes.
Code Format Scoring
N Scalar Rel. error $\leq$ 2%
B Boolean Exact match
S Sequence Fréchet $\leq$ 0.6 m
L List Set equality
C Coord. Spatial predicate
Table 4: Cognitive skill clusters.
Code Cognitive Target
M Metric computation
T Topology
G Geometric planning
V Visual logic / occlusion
R Rule-based validation
Structured layouts are flattened into lists of labeled objects, each defined by an axis-aligned bounding box. Object names are suffixed with instance identifiers (e.g., fridge_1, table_3) to ensure that referents are unique and stable across prompt construction and answer evaluation. All coordinates are expressed in meters, using a bottom-left origin and $(x,y)$ ordering. There is no hierarchical or typed schema beyond the flat object list, and no implicit conventions are assumed. A single system prompt is prepended to every interaction, specifying general and room-specific constraints: “You are a spatial assistant. Assume that all windows have zero thickness.” For layouts of type living room or bedroom, the prompt further clarifies that rugs may be overlapped by other objects, allowing for more realistic spatial arrangements involving soft surfaces.
Some question types are instantiated deterministically given the layout geometry (e.g., exact distances between centroids), while others are drawn from constrained samplers over feasible object pairs, free-space regions, or task-specific heuristics. The prompt templates themselves remain fixed across layouts and models. This uniformity allows us to isolate spatial reasoning under consistent linguistic and structural conditions.
These cognitive categories correspond to distinct reasoning faculties. Metric tasks (M) require explicit numerical computations over distances, areas, or positions, often grounded in real-valued geometric attributes. Topological tasks (T) depend on spatial relationships like adjacency or containment that are invariant under metric transformations. Geometric planning (G) involves multistep spatial inference over object configurations, including feasibility, collision avoidance, or spatial optimization (e.g., placing a large box without overlap). Visual logic (V) tasks such as visibility and angle estimation rely on implicit occlusion reasoning or directional projection. Rule-based tasks (R) involve symbolic constraints, such as verifying whether a layout satisfies clearance or adjacency requirements or whether essential objects are missing. These clusters are intended to reflect the diversity of tasks while enabling diagnostic grouping of error types. They also reflect a spectrum of abstraction, fromm low-level metric properties to high-level design affordances. While these clusters are not mutually exclusive, they provide a useful abstraction for analyzing model behavior and diagnosing failures. For example, we hypothesize that questions involving geometric planning and occlusion (G, V) will be disproportionately challenging for models without grounded spatial priors. This hypothesis is revisited in our results analysis (§4). The formatrmat and target codes are defined in Tables 4 and 4, and used throughout Table 2.
3.3 Evaluation Protocol and Scoring
Each question in PlanQA is paired with a reference answer derived entirely from the symbolic layout, allowing deterministic and automated evaluation of model outputs. Depending on the response type, correctness is assessed using tolerance-based numeric comparison, symbolic string matching, or geometric validation routines. The scoring logic referenced by each question format code is detailed in Table 4, and is applied uniformly across models.
For numerical questions (e.g., distances, areas, angles), predictions are accepted if they fall within a fixed relative error of the reference: $\left|\frac{\hat{y}-y^{*}}{y^{*}}\right|\leq\varepsilon$ where $\varepsilon$ is typically 0.02. This threshold is applied uniformly across models, and reflects spatial resolution limits in architectural layouts.
Binary and categorical responses, such as visibility, object fit, or violation detection, are validated against geometry-aware checks. For example, visibility is determined by ray intersection tests over the layout’s free space, while object fit questions are scored by verifying non-overlap and clearance constraints for a candidate placement. Boolean answers must match the reference exactly.
Structured outputs, including bounding boxes and paths, are checked for both format and semantic correctness. Rather than using exact coordinates, we validate these using geometric plausibility conditions—e.g., sufficient overlap with ground truth, or collision-freeness for planned paths. Thresholds for spatial alignment (e.g., bounding box proximity or path deviation) are set conservatively to account for minor lexical variation without over-penalizing near-correct reasoning.
To differentiate reasoning failures from generation or parsing errors, we annotate each model response with a status flag: Correct, Incorrect, Malformed, or Missing. Malformed responses include answers lacking the required structure “Final Answer:” or acceptable variations thereof, or containing ill-typed data. These annotations allow us to separate breakdowns in spatial inference from extraction failures and language model formatting idiosyncrasies.
3.4 Model Inference Setup
All models are queried using standard chat-based completion APIs, with prompts constructed as described above. We evaluate a mixture of open-weight and proprietary language models across three resource tiers, distinguished by their token limits and relative inference cost. Models in the “reasoning” tier are allocated up to 8192 tokens per completion, enabling them to process longer layout descriptions and multi-step CoT outputs. Other models operate within a 4096-token context window and are evaluated only on the non-reasoning subset of questions.
No model receives fine-tuning or prompt adaptation specific to PlanQA. All prompts are zero-shot, with the system message and output formatting constraints held fixed. For each model, we use the default inference configuration provided by its vendor, including temperature settings. In practice, these are near-deterministic across runs and sufficiently stable for head-to-head comparison. In cases where a model produces variable-length completions or truncates early, we retain the first parseable instance of the “Final Answer:” line, if present.
Each model is evaluated over a batch of 600 layouts per room type, with one question per layout. Requests are grouped into new-line-delimited JSON (JSONL) payloads for efficient submission via batched APIs. Layout serialization, prompt insertion, and output parsing are fully automated, with no manual curation or intervention.
We include both high-capacity generalist models and compact instruction-tuned variants to measure the degree to which spatial reasoning emerges with scale and supervision alone. All models are evaluated on identical input distributions and scoring criteria, enabling cross-system comparisons that are both cost-aware and architecture-agnostic.
4 Results
We evaluated the performance of the model on a dataset of 1,800 layouts, consisting of 600 kitchens, 600 living rooms and 600 bedrooms. Each layout is paired with one question sampled from a pool of 10 parameterized templates, filtered by room applicability. This yields 18,000 layout–question pairs. The full taxonomy is shown in Table 2.
Eleven language models are evaluated, covering a wide range of parameter scales, architectures, and training regimes. The reasoning-oriented models include Qwen3-32B Yang et al. (2025) and DeepSeek-R1 DeepSeek-AI (2025). The large general-purpose models include DeepSeek-V3 DeepSeek-AI (2024), LLaMA 3.3-70B-Instruct, Gemma 2-27B-it Riviere et al. (2024), phi-4 Abdin et al. (2024b), and GPT-4.1 OpenAI et al. (2024). The smaller models include LLaMA 3.1-8B-Instruct Grattafiori et al. (2024), Gemma 2-9B-it Riviere et al. (2024), Phi-3.5-mini Abdin et al. (2024a), and GPT-4o-mini. All models are evaluated in identical zero-shot conditions using standardized prompts and serialized layout inputs, as described in Section 3. The complete results disaggregated by question, room type, and model are included in the appendix.
4.1 Quantitative Results
We begin by aggregating accuracy across cognitive targets as defined in Table 4. Figure 2 shows a heatmap of mean accuracy by model and target category. Metric reasoning (M) and visual logic (V) show the highest cross-model scores, while Topology (T), rule-based validation (R), and geometric planning (G) are more difficult across the board. DeepSeek-R1 (reasoning), GPT-4.1 (large), and GPT-4o-mini (compact, tuned) consistently lead the performance in the harder categories, although even large models struggle in certain occlusion and clearance tasks. Surprisingly, GPT-4.1’s performance on vector-direction (V) questions nearly matches DeepSeek-R1 and GPT-4o-mini—despite those models’ explicit chain-of-thought optimization—highlighting an unexpected parity in occlusion reasoning.
Figure 2: Accuracy by model and cognitive target. Each cell shows average accuracy across questions associated with a given skill cluster. Cognitive targets: M = Metric computation, T = Topology, G = Geometric planning, V = Visual logic / occlusion, R = Rule-based validation.
Performance improves with model scale, but not uniformly across tasks. As shown in Figure 3, accuracy improves steeply up to  30B parameters, suggesting that scale alone cannot fully close the gap on complex spatial tasks. Notably, gains largely plateau beyond 30 B parameters, indicating diminishing returns for raw scale on planning and constraint problems.
Figure 3: Accuracy vs. model size. Models are arranged from the smallest to largest in terms of trainable parameters. While larger models perform better overall, instruction tuning and task specialization also contribute significantly.
We also compare accuracy across room types (Figure 4). Kitchens exhibit the highest overall accuracy, likely because their regular counter-and-appliance grids reduce occlusion and simplify clearance reasoning. In contrast, living rooms and bedrooms generally yield lower performance across models, due to denser layouts, complex furniture adjacencies, and higher occlusion likelihood, with frequent object groupings such as televisions on stands, sofas and coffee tables on rugs, or lamps on nightstands. These room-type gaps suggest that layouts with strong grid structure provide easier inductive priors for LLMs than organic, cluttered arrangements.
Figure 4: Accuracy by question and room type. Kitchens show the highest average accuracy, likely due to their regular, grid-like counter layouts and lower occlusion. The inability of most models to estimate free space in living rooms and bedrooms may be caused by more clutter, or the presence of objects like area rugs which the model was instructed to ignores for occlusion purposes.
4.2 Ablation
To evaluate the robustness of layout interpretation under alternate encodings, we perform a format ablation that replaces the standard JSON layout with a semantically equivalent XML version. This substitution preserves all geometric and object-level content while modifying only the syntax and structural serialization. To reduce evaluation cost, we focused on the most variance-sensitive question categories in the kitchen domain: Repositioning, Vector Direction, and Free Space. These were selected based on standard deviation in accuracy across the full model set. We apply each ablation to three illustrative models spanning architecture classes and capacity tiers: GPT-4.1 (closed large), GPT-4o-mini (compact tuned), and DeepSeek-R1 (reasoning-optimized). As shown in Table 5, accuracy remains stable across formats for most question types, suggesting that these models encode layout semantics in a manner invariant to low-level representation details.
Table 5: Accuracy using JSON vs XML layout encoding. Each cell shows performance on the original JSON representation and its equivalent XML rendering.
Model Repositioning View Angle Obstruction
GPT-4.1 77.0 → 76.5 92.2 → 93.7 75.5 → 89.6
GPT-4o-mini 33.8 → 29.6 42.4 → 42.2 12.2 → 15.8
DeepSeek-R1 83.6 → 93.7 74.7 → 74.3 95.7 → 96.1
5 Conclusion
We have presented PlanQA, a benchmark that evaluates spatial reasoning over symbolic 2D layouts by leveraging object-centric encodings (JSON, XML) similar to those used in architecture and robotics. Through a systematic evaluation of 11 language models across five skill clusters: metric computation, topology, geometric planning, visual logic, and rule-based validation, we make critical findings with implications for the use of LLMs in geometric design tasks.
Metric and basic visibility queries (distances, areas, line-of-sight) are handled reliably (>95% accuracy), whereas geometric planning, topological relations, and constraint verification consistently fall below 50% accuracy. Performance gains level off beyond approximately 30 billion parameters, and models optimized for chain-of-thought show only modest improvements on the hardest tasks. Additionally, grid-like layouts (kitchens) yield higher accuracy than more varied configurations (living rooms, bedrooms), indicating that layout regularity influences model performance.
These results reveal that current LLMs do not form sufficiently robust internal representations for complex spatial inference. Addressing this limitation could take two complementary directions. In the near term, integrating external geometric solvers or symbolic planning modules can compensate for the models’ weaknesses in collision avoidance and clearance reasoning. In the longer term, training regimes that incorporate explicit spatial constraints—through specialized objectives, data augmentation with irregular layouts, or constraint-violation examples—may be necessary to improve performance on design-oriented tasks.
Acknowledgments
The work is supported by funding from KAUST, Center of Excellence for Generative AI, under award number 5940, and the NTGC-AI program
Appendix A Technical Appendices and Supplementary Material for PlanQA
This technical appendix includes an overview of the dataset generation process, followed by full evaluation results on our generated QA pairs, along with aggregated metrics and ablation studies. It also presents an extended taxonomy of spatial reasoning tasks, explanations of selected failure cases, an expanded analysis of vision-language models (VLMs), and examples of prompts used for both generation and questioning.
The full dataset and all code used for generation, evaluation, and visualization are provided in the supplementary materials.
Appendix B Details on Synthetic Data Generation
Room layouts were generated using Gemini 2.5 Pro, specifically fine-tuned for spatial reasoning and bounding-box tasks Team et al. (2025). Below, we describe the detailed procedure and constraints. All generation scripts, prompts, constraint-checking code, and random seeds are included in the code for full reproducibility.
B.1 Room Shape Generation
We generated 600 layouts each for kitchens, bedrooms, and living rooms. These layouts include a variety of geometries and sizes, with clearly defined room structures containing windows, doors, and corner cutouts where applicable. All layouts follow specifications tailored to each room type.
Each layout falls into one of three shape categories: rectangular, L-shaped, or open. Size categories were defined individually for each room type and were incorporated into the generation prompts based on common assumptions about typical room dimensions.
The distribution of room shapes and sizes is shown in Table 6, and example layout types are illustrated in Fig. 1.
Table 6: Shape and size distribution across generated layouts for each room type.
Room Type Shape Distribution (Rect / L-shaped / Open) Size Categories (in m2) (Small / Medium / Large)
Kitchen 40% / 40% / 20% $\leq$7 / 7–18 / $>$18
Bedroom 50% / 30% / 20% 8–12 / 12–18 / $>$18
Living Room 40% / 40% / 20% 20 / 22 / 24
Geometric and Structural Constraints:
Room geometries were procedurally varied using prompt-based guidance to produce rectangular, L-shaped, and open-plan configurations. The following structural properties were described in the prompts but not explicitly enforced during layout generation:
  • L-shaped rooms were described as rectangular spaces with a square cutout from one corner. To maintain usable proportions, each leg of the L shape was suggested to be at least 1.5,m wide and deep.
  • Open-plan rooms, apart from living rooms, were prompted without doors and with one full wall removed. This was intended to vary room shapes and simplify layout generation.
  • Doors were described with widths between 0.8,m and 1.0,m, selected randomly. Windows were chosen from a fixed set of widths: 0.6,m, 0.75,m, 0.9,m, 1.2,m, and 1.5,m.
  • Prompts included guidance to place all elements such as doors, windows, and cutouts entirely within room boundaries.
Window Placement:
Window placement followed prompt-based guidelines aimed at supporting daylight access and layout clarity:
  • The total window length was set to exceed 15 percent of the room’s floor area. It was used as a simple proxy to ensure visible window openings and visual balance along the walls.
  • Windows on the same wall were the same size to support visual balance. Small, isolated windows were not used.
  • Long windows, over 1.5 m, were split into segments with 0.05 to 0.15 m gaps for a more modular appearance.
  • Windows were not placed opposite each other or on the same wall as a door, as such arrangements are less common and were not emphasized in the prompt design.
B.2 Furniture and Appliance Placement
In the second stage, each room was populated with furniture and appliances based on layout style and object-specific constraints. While the styles differ by room type–such as enforcing a work triangle in kitchens, orienting seating around a focal point in living rooms, or centering beds symmetrically in bedrooms–the overall placement process followed a unified set of rules:
  • All floor-standing objects must be placed without overlaps, except in semantically grouped cases (e.g., lamps on tables, chairs under tables).
  • Clearance zones must be preserved around doors, main pathways, and functional elements such as beds, appliances, and desks.
  • Placement follows a priority order: essential furniture is placed first, followed by optional and decorative elements only if space allows.
  • Major objects like fridges, ovens, and beds must be anchored to structural walls or room boundaries.
Layouts violating any of these hard constraints, due to overlap, clearance issues, or improper attachment, were automatically discarded.
B.3 Layout Selection Criteria
Approximately one-third of the initially generated room layouts were filtered-out using a set of geometric and functional constraints. These filters were designed to ensure realistic object placement, functional usability, and architectural plausibility. While these rules are based on general design principles, they are not based on any specific design standard. Instead, they are the result of iterative development, focusing specifically on addressing cases where our prompts lead to unlikely or implausible layouts. After filtering, we retained 600 valid layouts for each room type.
  • Non-overlapping objects (with exceptions): Each pair of objects must satisfy axis-aligned bounding box (AABB) separation constraints, unless they belong to a known exception category. For two objects $A$ and $B$ with bounding boxes $(x_{1}^{A},y_{1}^{A},x_{2}^{A},y_{2}^{A})$ and $(x_{1}^{B},y_{1}^{B},x_{2}^{B},y_{2}^{B})$, non-overlapping requires:
    $x_{2}^{A}\leq x_{1}^{B}\quad\text{or}\quad x_{2}^{B}\leq x_{1}^{A}\quad\text{or}\quad y_{2}^{A}\leq y_{1}^{B}\quad\text{or}\quad y_{2}^{B}\leq y_{1}^{A}$
    This constraint is not enforced for the following semantically compatible object pairs: (i) rug with any object placed on top of it; (ii) lamp with nightstand, desk, or table; (iii) tv with tv_stand; (iv) chair objects with desk or table.
    These exception pairs are considered contextually collocated or hierarchically related (e.g., support/surface relationships) and are therefore allowed to overlap.
  • Non-blocking door clearance: Doors have physical thickness and are defined by bounding boxes $(x_{1}^{d},y_{1}^{d},x_{2}^{d},y_{2}^{d})$. A clearance zone of $door\_length$ meters is required in front of the door to ensure swing space and accessibility. The position of this zone depends on which wall the door is attached to.
  • No windows on opposite walls: We did not include layouts with windows on directly opposite walls, as such configurations are uncommon in typical residential designs.
  • Appliances against walls or cutout edges: Large fixtures like fridges and ovens must be flush against at least one wall or cutout boundary. This is formalized by enforcing:
    $x_{1}=0\quad\text{or}\quad x_{2}=W\quad\text{or}\quad y_{1}=0\quad\text{or}\quad y_{2}=D\quad\text{or edge of cutout}$
    For rooms with cutouts, an object may align with a cutout boundary, defined as additional wall segments with known coordinates.
These constraints were iteratively selected to address common implausible layouts generated by Gemini 2.5 Pro using our prompts. They are not universal requirements for real layouts, nor a complete set of constraints, but aim to avoid frequent sources of implausibility in generated layouts.
B.4 Room Statistics
Table 7 summarizes key statistics for the generated layouts across room types, including average object count, overlap frequency, and spatial density. Kitchens typically contain fewer objects but have higher density due to their compact size. In contrast, bedrooms and living rooms host more objects with greater overlap, reflecting more open layouts and flexible furnishing.
This structural simplicity, with fewer objects, minimal overlap, and higher density, makes kitchens the simplest environment for spatial reasoning tasks. Due to the limited amount of free space, some complex questions, such as those involving paths, have fewer possible answers. As a result, models tend to perform better on kitchen-related questions.
Table 7: Average layout statistics by room type.
Metric Kitchen Bedroom Living Room
Avg. # of Objects 13.30 14.24 15.29
Avg. # of Overlaps 0.52 6.27 6.26
Avg. Object Density 1.15 0.84 0.71
Appendix C Full Breakdown by Room Type, Question Category, and Model
Table 8 reports per-model accuracy separated by room type (kitchen, living room, bedroom) and question category. This is the full unaggregated results matrix underlying the analyses presented in Section 4. The table includes all eleven evaluated models and eleven question types. Path is assessed using two complementary metrics: validity and Fréchet Distance. Together, these question types span a range of challenges, covering both reasoning-heavy and functionally grounded tasks.
The table presents accuracy for all eleven models across eleven question types, grouped by room type. Each room type–kitchen, bedroom, and living room–includes results for all question types, totaling 33 rows. Each value reflects model accuracy over a fixed set of 600 questions, corresponding to 600 layouts per room type. For every combination of question type and room type, the same set of 600 questions is used across all models to ensure fair comparison.
Table 8: Question-level accuracy by category for each model across room types. Models (left to right): Qwen3-32B, DeepSeek-R1, DeepSeek-V3, LLaMA 3.3-70B-Instruct, Gemma 2-27B-it, Phi-4, GPT-4.1, LLaMA 3.1-8B-Instruct, Gemma 2-9B-it, Phi 3.5-mini, GPT-4o-mini.
Note: High truncation rate due to token limits.
Reasoning Big Models Small Models
N 32.8B 671B 671B 70B 27.2B 14B N/A 8B 9.2B 3.8B N/A
Temp. 0.6 0.6 0.3 0.0 0.5 0.0 1.0 0.6 0.5 0.7 1.0
Type Category [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
K Distance 97.2 99.5 98.5 97.2 94.7 90.7 100 82.3 74.2 35.3 97.3
Area (counters) 85.3 99.5 95.7 81.7 46.5 79.2 98.5 30.8 37.5 9.8 80.3
Free Space 83.7 88.0 52.5 37.0 23.5 42.0 83.5 9.5 13.5 6.3 20.3
View Angle 74.0 58.2 69.2 75.2 12.7 52.8 86.3 11.5 9.2 8.2 42.2
Repositioning 91.5 96.3 69.8 48.7 14.8 22.7 85.8 3.7 6.2 11.5 40.7
Max Box 33.3 30.7 9.2 8.8 3.0 3.0 57.2 1.7 1.2 1.7 3.7
Fit/Placement 92.8 92.5 71.0 66.2 72.0 82.0 89.8 68.7 71.8 68.8 70.7
Path (Valid) 13.2 6.3 39.3 30.3 23.2 26.8 73.0 10.5 17.5 14.8 35.7
Path (Fréchet) 15.0 6.3 32.8 30.2 26.0 26.0 55.8 10.2 19.0 9.3 36.7
Missing Object 87.3 88.7 44.3 56.2 52.3 52.2 79.8 27.5 39.7 14.3 58.0
Obstruction 84.0 95.2 32.7 6.0 3.3 9.7 93.5 1.2 2.7 11.3 14.2
L Distance 98.7 99.8 99.5 98.2 96.3 98.8 99.8 87.0 81.8 58.8 98.5
Area (sitting) 96.7 99.5 84.5 98.0 83.8 88.7 99.5 32.5 41.3 12.8 85.7
Free Space 0.3 4.3 1.0 0.3 3.7 0.3 5.0 0.7 3.2 1.5 1.7
View Angle 81.5 86.0 70.0 76.8 14.3 50.3 96.0 14.8 8.3 10.5 42.2
Repositioning 80.5 93.0 45.0 33.5 12.8 19.2 71.5 6.5 5.7 6.3 29.3
Max Box 1.5 0.8 1.0 3.0 2.0 2.8 6.5 0.8 1.5 1.0 2.3
Fit/Placement 90.7 91.2 71.5 80.0 83.7 87.7 91.8 75.0 75.5 72.8 72.7
Path (Valid) 10.0 14.7 33.0 26.7 30.5 26.7 53.2 7.3 21.7 16.7 33.5
Path (Fréchet) 13.2 14.2 23.7 17.8 19.2 14.7 48.0 2.5 14.2 6.8 25.2
Missing Object 73.0 76.2 51.0 49.3 29.5 36.0 65.5 9.7 28.3 11.7 32.3
Obstruction 80.7 96.5 24.3 7.3 3.8 9.3 84.7 2.5 5.2 4.7 11.7
B Distance 98.7 99.8 99.5 98.2 96.3 98.8 99.8 87.0 81.8 58.8 98.5
Area (storage) 98.7 99.8 94.0 97.0 86.3 88.3 99.3 30.3 66.3 47.7 88.2
Free Space 1.7 5.8 1.2 0.3 2.5 1.2 2.8 1.8 1.0 1.0 1.2
View Angle 76.0 79.8 70.0 78.3 10.8 57.0 94.2 15.3 10.2 7.7 43.0
Repositioning 78.7 94.3 53.8 36.0 11.2 15.7 73.7 4.3 9.0 6.5 31.5
Max Box 0.7 1.0 2.0 1.8 2.0 2.5 7.2 1.0 0.7 1.0 2.0
Fit/Placement 86.3 86.3 65.8 66.7 66.5 73.0 82.0 66.0 63.8 64.7 66.2
Path (Valid) 15.5 20.8 49.7 40.7 38.5 35.8 67.8 11.3 34.5 20.5 48.8
Path (Fréchet) 15.3 21.5 30.3 30.3 27.0 19.0 49.5 3.7 19.7 8.5 36.3
Missing Object 64.3 65.2 39.8 40.3 33.0 25.2 61.0 14.8 19.7 8.3 34.5
Obstruction 87.2 95.7 32.0 4.3 1.3 5.8 89.5 2.2 5.3 9.3 12.2
Appendix D Additional Ablations and Analysis
To further verify the robustness and generalizability of our evaluation setup, we conducted several supplementary analyses beyond the primary results in Table 8.
Mean and Variability Across Room Types
Figure 6 shows radar plots summarizing model performance across question types, averaged by room type. Kitchens, living rooms, and bedrooms demonstrate similar high-level trends: models are generally most accurate on questions involving spatial metrics such as Distance, Area, and Fit/Placement, while struggling more with abstract or multi-object tasks such as Max Box and Path (Valid).
Figure 6 illustrates the standard deviation across models per question category. Notably, Repositioning, View Angle, Area, and Obstruction show the highest variability across models, particularly in bedrooms and living rooms. Free Space, in contrast, exhibits high variance only in kitchens, highlighting that performance inconsistency is not uniform across room types.
Figures 8 and 8 provide complementary radar plots organized by model, highlighting differences in room-specific performance. Kitchens consistently exhibit slightly higher mean accuracy across models, particularly for larger models like GPT-4.1 and DeepSeek-R1. Variance across models is relatively uniform, though larger models tend to show reduced variance overall.
Figure 5: Mean accuracy by question type across room categories. Each axis corresponds to a specific spatial reasoning skill (e.g., distance, placement, path validity).
Figure 6: Standard deviation in model accuracy across question types and room categories. Higher values indicate greater variability across models within a room type.
Figure 7: Mean accuracy across models by room category. Shaded regions reflect performance in bedrooms, kitchens, and living rooms, highlighting which models generalize best across spatial tasks.
Figure 8: Standard deviation of model performance across room types. Captures how consistent each model is across different spatial environments. Narrower shapes indicate more uniform behavior.
Prompt Sensitivity Ablation
To assess model robustness to prompt variation, we conducted an ablation in which each question was regenerated using the same template but with alternate object references. For example, a question originally referring to a "sofa" might instead use a "bookshelf" in the regenerated version. This allowed us to evaluate whether model performance remains stable under changes in object content while preserving linguistic structure.
As shown in Table 9, performance generally remained consistent, with only minor differences across models. Larger models such as GPT-4.1 were more stable, while smaller models like GPT-4o-mini showed slightly higher sensitivity. These results suggest that the evaluation framework is robust to this type of variation.
Table 9: Accuracy under prompt variation. Each cell shows performance on the original prompt followed by a regenerated version with alternate object references.
Model Repositioning View Angle Obstruction
GPT-4.1 77.0 → 75.3 92.2 → 91.4 75.5 → 74.9
GPT-4o-mini 33.8 → 32.9 42.4 → 46.7 12.2 → 14.1
DeepSeek-R1 83.6 → 95.7 74.7 → 72.1 95.7 → 95.6
Layout Perturbation Ablation
To test model sensitivity to small physical changes, we introduced random spatial jitter of up to ±0.1 m to all object coordinates during evaluation. This simulated minor but plausible noise in the layout. After jittering, we recomputed the correct answers for all QA pairs based on the updated room configuration. The questions remained unchanged, but all evaluations used the perturbed layouts.
The jitter was constrained to ensure that objects did not overlap, remained within room boundaries, and did not block doors. Jittering can affect spatial relationships such as clearances, alignments, and angles. For example, a path that was previously clear might become blocked, or an object may no longer fit within a gap.
As shown in Table 10, model accuracy showed slight fluctuations in response to the spatial changes. Overall, model performance remained robust, indicating resilience to small, physically realistic layout variations.
Table 10: Accuracy before and after layout jitter. Values shown for three spatially sensitive questions.
Model Repositioning View Angle Obstruction
GPT-4.1 77.0 → 72.5 92.2 → 87.3 75.5 → 80.2
GPT-4o-mini 33.8 → 28.4 42.4 → 42.5 12.2 → 11.8
DeepSeek-R1 83.6 → 89.8 74.7 → 70.3 95.7 → 97.1
Appendix E Spatial Reasoning Question Suite
To extend our taxonomy of question types introduced in Table 2, Table 11 details the specific evaluation procedures used to compute ground truth answers for each category. In addition to categorizing the task, the table shows how answers are derived and what types of spatial reasoning each task demands. This breakdown clarifies not only what is being tested, but how it is measured, supporting both reproducibility and deeper analysis of model performance in later sections.
Table 11: Categories of spatial reasoning questions, answer-checking procedures, and diagnostic insights.
Category Sample Question Answer Checking Diagnostic Insight
Distance What is the Euclidean distance between the center of the door and the fridge? Compute centroids $(x,y)$ of both boxes; calculate Euclidean distance between them; compare with model output allowing a tolerance of $2\%$ of the ground-truth distance. Tests numeric computation, coordinate extraction, and arithmetic precision.
Area What is the area of the kitchen counter? Multiply the width $\times$ height of each cabinet; sum all resulting areas; compare the total against the model output with a 2% tolerance. Probes multiplicative reasoning, aggregation, and unit consistency.
Free Space What is the total unoccupied floor area in the bedroom? Merge overlapping object bounding boxes; sum the resulting areas; subtract from total room area; compare the unoccupied area with model output using a 2% tolerance. Measures the ability to aggregate and subtract multiple regions.
View Angle What is the dot product between the facing vector of the sofa and the vector to the TV? Compute normalized facing vector of reference; get normalized vector to target center; compare dot product with model output using 2% tolerance. Evaluates directional reasoning, vector normalization, and spatial alignment understanding.
Repositioning How far can the sofa move right before hitting a wall or another object? Simulate axis-aligned movement; measure distance to furthest valid non-overlapping position; compare with model output using 2% tolerance. Tests spatial simulation, collision detection, and directional constraint reasoning.
Max Box What is the largest axis-aligned box you can place in the living room? Use brute-force to solve largest empty rectangle in free space; compare resulting dimensions with 2% tolerance. Probes continuous-space optimization and spatial coverage estimation.
Fit/Placement Can a $2\,\mathrm{m}\times 3\,\mathrm{m}$ table fit in the room without overlap? Attempt to place a rectangular object in free space; ensure no intersection with existing bounding boxes. Compare boolean result with model prediction. Tests collision detection, spatial constraint enforcement, and free-space reasoning.
Shortest Path What is the shortest walkable path from the fridge to the exit with $15\,\mathrm{cm}$ clearance? Offset free space by $15\,\mathrm{cm}$ for clearance; use A* algorithm on navigable grid to compute path; compute Fréchet distance (threshold $0.6\,\mathrm{m}$) to model path; check its validity. Tests path planning, clearance-aware modeling, and heuristic graph search.
Missing Object Where should the missing armchair be placed in this layout? Check if predicted bounding box fits in free space without overlaps and satisfies clearance/accessibility constraints. Evaluates spatial reasoning, free-space utilization, and functionally appropriate object placement.
Obstruction Which objects obstruct the line of sight between the bed and the window? Compute vector from target to reference center; check which bounding boxes intersect the segment (excluding endpoints). Tests geometric reasoning, line–rectangle intersection, and occlusion detection.
Appendix F Failure Cases
To better understand the sources of model failure, we conducted a qualitative analysis of representative examples from the benchmark. This section presents visualizations of selected test layouts alongside model responses. By examining both correct and incorrect outputs, we aim to identify common failure patterns and reasoning bottlenecks across different architectures.
F.1 Max Box
As illustrated in Figure 10, this example shows GPT-4.1 producing a suboptimal result in a spatial reasoning task. Instead of identifying the largest free region in the kitchen layout, the model placed a bounding box that extended along a single axis, starting from one edge, without fully exploring the layout. It selected an area from x=0 to 2.1 and y=1.1 to 1.8, yielding 1.47 m² of free space.
A larger available region, shown in Figure 10, spans from (0.7, 0.7) to (2.1, 1.8), and offers 1.54 m². The model did not identify this larger region, which indicates that its search strategy was likely not exhaustive, even though it had access to the full structured input. The reasoning trace in Figure 11 confirms that it evaluated a set of candidate regions but ultimately settled on a suboptimal one. The criteria for candidate selection remain unclear and may vary across queries.
Figure 9: GPT-4.1 prediction
Figure 10: Ground truth
LLM Reasoning Summary Let’s analyze the problem step by step. First, I’ll look for large free areas without obstacles. Option 1: The region to the right of the chairs and table, from x=1.3 to 2.1, and from y=0 to 1.8. This gives an estimated area of 1.44 m². Option 2: The space above the chairs and below the wall units, from x=0 to 2.1 and y=1.1 to 1.8. This gives an area of 1.47 m². Option 3: A small gap near the bottom right corner. It looks too narrow and likely blocked by furniture. Option 4: The space in front of chair_2, but it appears too small to be useful. Based on these, Option 2 seems best. I’ll choose the rectangle [0.0, 1.1, 2.1, 1.8], which gives an area of 1.47 m².
Figure 11: Summarization of chain-of-thought style output from GPT-4.1. The model imitates reasoning by listing and evaluating candidate regions, then selects a suboptimal solution.
F.2 Shortest Path
In this task, the model must find a path from the center of a starting object to the center of a destination object. For simplicity, the start and goal objects are not treated as obstacles.
Our pathfinding system constructs a visibility graph by connecting all mutually visible corner points in the environment. Dijkstra’s algorithm is then used to compute the true shortest path over this graph, ensuring optimality under Euclidean distance. This method avoids the limitations of grid-based approximations and produces precise, flexible paths around obstacles. While multiple valid solutions may exist depending on object layout and clearance constraints, the chosen path always reflects the shortest feasible route.
When interpreting the task, models often fail to make effective use of available free space. Instead of navigating around obstacles or adapting to the environment, they frequently extend a straight line in one direction until it is blocked. This leads to incomplete or inefficient trajectories that do not reflect the geometry of the room or required clearance. In some cases, paths cut too closely around objects, likely due to the model implicitly minimizing Euclidean distance without accounting for spatial constraints.
Figure 13 shows a path that reaches the target but fails the Fréchet distant constraint of 0.6 m, making it non-optimal. Figure 13 depicts a path that violates clearance constraints and does not represent a valid trajectory to the target.
Figure 12: Comparison between the optimal path (green) and the GPT-4.1-generated trajectory (cyan), which, despite reaching the target, exceeds the Fréchet distance threshold of 0.6 m.
Figure 13: Depiction of the optimal path (green) alongside the GPT-4.1-generated trajectory (red), which fails to satisfy clearance requirements and results in an infeasible route.
Appendix G VLM Experiments
To further probe model performance on particularly challenging tasks, we conducted targeted evaluations of two vision-language models (VLMs). We focused on the most difficult question categories–Free Space and Max Box–within the most spatially complex environments: living rooms and bedrooms. The models tested in this focused benchmark include Gemini 2.5 Flash Preview 04-17 Gemini Team, Google (2025) and GPT-4.1. Results are summarized in Table 12, which compares each model’s output under two input conditions: JSON only and JSON combined with a floorplan image. An example of Gemini’s step-by-step reasoning under the JSON+Image condition is shown in Figure 14.
Table 12: Comparison of spatial understanding between Gemini 2.5 Flash and GPT-4.1 for Living Room (L) and Bedroom (B), under different conditions.
Type Category [Uncaptioned image] Gemini 2.5 Flash [Uncaptioned image] GPT-4.1
JSON Only JSON + Image JSON Only JSON + Image
L Free Space 6.4 12.2 5.0 9.8
Max Box 17.0 22.0 6.5 4.3
B Free Space 12.8 15.2 2.8 10.5
Max Box 15.7 21.0 7.2 7.2
[…]
Figure 14: Summarized step-by-step reasoning by Gemini to compute maximum free space from a JSON floorplan and image. The model merges nested objects and analyzes rectangular gaps to estimate usable area.
While our results show that both models benefit from visual input in some settings, the extent of improvement varies by task. For the Free Space category, both Gemini 2.5 Flash and GPT-4.1 show substantial gains when provided with floorplan images. In contrast, for Max Box, only Gemini shows a notable improvement, while GPT-4.1 performs similarly or worse with visual input. This variation suggests that visual grounding supports some spatial tasks more effectively than others.
The stronger visual performance of Gemini 2.5 Flash may stem from architectural or training differences, possibly reflecting design choices aimed at robotics or spatial reasoning tasks.
One challenge lies in the nature of floorplan inputs. Unlike natural images, floorplans vary widely in style, abstraction, and labeling. This inconsistency makes it harder for models to extract spatial meaning reliably.
Smaller models show little benefit from visual input, likely due to limited capacity and lack of pretraining for structured spatial tasks.
In summary, floorplan understanding requires more than object detection. It demands precise spatial reasoning and integration with structured formats like JSON. Without task-specific tuning or standardized visual formats, even advanced models remain limited in this domain.
Appendix H Prompts
This section presents selected prompt templates used in our benchmark. We include a summarized version of the bedroom layout generation prompt, as well as representative examples of question prompts targeting specific spatial reasoning skills. These prompts reflect the structured task design outlined in Table 11. Additional prompts–including full versions, formatting logic, and implementation details–are provided in the supplementary code and materials for reproducibility.
Prompt: Max Box You’ll be given a layout of a {room_type} in JSON format. Your task is to calculate the area (in m²) of the largest axis-aligned rectangle that can fit within the room. The rectangle must not overlap with any other objects, but it may fully occupy free space, including areas above rugs. You can ignore passageways, door clearance, or accessibility constraints. Room layout: {room_json} Reason step by step to calculate the area. Respond only with a float value in square meters using the format: *Final answer*: area
Figure 15: Prompt for computing the largest empty rectangle area within a room layout using Chain-of-Thought reasoning.
Prompt: Missing Object Placement You’ll be given a layout of a {room_type} in JSON format. Your task is to suggest a suitable placement for the missing object {object_name} by providing its bounding box coordinates. The placement must avoid overlaps, respect obstacles, and maintain accessibility. Room layout: {room_json} Reason step by step to identify the placement. Respond only with bounding box coordinates (float values) using the format: *Final answer*: [x_min, y_min, x_max, y_max]
Figure 16: Prompt for placing a missing object in a room layout using Chain-of-Thought reasoning. The model must return valid coordinates that satisfy geometric constraints and maintain functional space.
[…]
Figure 17: Summarized data generation prompt for producing structured and constrained bedroom layouts.
Prompt: Fill Bedroom Layout with Objects Given a predefined bedroom layout style and room geometry in JSON format, generate a filled 2D bird-view layout. Include a list of placed objects with their bounding boxes and explanatory comments. Essential fields: Each object must have a "label", "bbox" ([y0, x0, y1, x1]), and a descriptive "comment". Furniture labels include: "bed", "nightstand", "dresser", "wardrobe", "desk", "chair", "armchair", "rug", "lamp", etc. Architectural elements ("door", "window", "cutout_area", "fireplace", "closet_alcove") must match the input layout and remain unmodified. Placement priorities: 1. Place the "bed" according to the bed_size_suggestion and layout style. 2. Add essential storage: "dresser", "wardrobe", or use "closet_alcove" if defined. 3. Add secondary items (e.g., "nightstand", "desk", "chair") only if space and clearance allow. 4. Add decorative or optional items ("rug", "mirror", "floor_lamp", "plant") last. Constraints: Maintain at least 0.75 m clearance for walkways and door swing. Beds require 0.6–0.75 m of access space on sides and foot (unless against wall). Wardrobes/dressers need 0.6–0.8 m clearance for drawer/door use. No object overlap (except table lamps on nightstands or rugs under furniture). Use walls efficiently; avoid blocking windows unless unavoidable. Ensure mirror has 0.75 m clearance in front; treat "rug" as an anchor but optional. Final output: a JSON list of objects, including placement and comments. No layout geometry should be altered.
Figure 18: Prompt for populating a bedroom layout with functionally and spatially valid object placements, following layout-specific design rules.